OpenStudy (anonymous):

1/m-2=2/m^2-6m+8-m^2+m-2/m2-6m+8 **Open this to find a much more neater version of this problem

OpenStudy (anonymous):

\[\frac{ 1 }{ m-2 }=\frac{ 2 }{ m^2-6m+8}-\frac{ m^2+m-2 }{ m^2-6m+8 }\]

OpenStudy (anonymous):

FYI: I multiplied both sides by m^2-6m+8 to try and cancel out the denominator or make a one out of it, but that didn't work.

OpenStudy (anonymous):

@agent0smith Can you help me with this?

OpenStudy (agent0smith):

Factor everything, is always a good idea in these.

OpenStudy (anonymous):

\[\frac{ 1 }{ m-2 }=\frac{ 2 }{ (m-4) (m-2)}-\frac{ (m+2)(m-1) }{ (m-4)(m-2) }\]

OpenStudy (anonymous):

@agent0smith that's what I got when I factored it.